OpenStudy (anonymous):

find the equation of the tangent line to the curve y+sin(x+2y)=1 at the origin

OpenStudy (cwrw238):

differentiate dy/dx + cos(x + 2y)* (1 + 2 dy/dx) = 0 now find dy/dx in terms of x and y to give the slope of the tangent line then use the general form y - y1 = m(x - x1) where (x1,y1) will be (0,0)

OpenStudy (anonymous):

ok sir thanks for the idea :) i'll just follow it to answer my homework :)

OpenStudy (cwrw238):

oh: - when you have done the first part (finding the slope m) plug in x = 0 and y = 0